# Impurity bands, line-nodes, and anomalous thermal Hall effect in Weyl superconductors

### v1

@@ IV Quasiclassical transport theory
IV.1 Eilenberger equation
The quasiclassical transport (Eilenberger) theory describes superconductors in the limit $1/(k_{\rm F}\xi_{0})\sim T_{\rm c}/T_{\rm F}\ll 1$. In this regime, the normal state DOS can be taken to be energy-independent over the range where the Gor’kov Green’s function is peaked [77]. Integrating this Green function over the band kinetic energy, $\xi_{\bm{k}}={\bm{k}}^{2}/2m-\epsilon_{\rm F}$, we define the quasiclassical Green function as,
@@ IV.1 Eilenberger equation
The quasiclassical Green function is defined at the Fermi surface and is the central object of the quasiclassical transport theory [78]. The superscript $\rm X=R,A,K$ in Eq. (9) represents the retarded, advanced, and Keldysh matrix elements, respectively. $\check{\tau}_{i}\;(i=x,y,z)$ are the Pauli matrices in the Nambu (particle-hole) space. Throughout this paper, we denote $\check{A}$ as a $8\times 8$ matrix in the Keldysh space and $\underline{A}$ as a $4\times 4$ Nambu matrix. If a matrix $\underline{A}\;(\check{A})$ is defined in the Nambu (Keldysh) space, the corresponding matrix $\check{A}\;(\underline{A})$ in the Keldysh (Nambu) space is defined as $𝟙\check{A}=\underline{A}\otimes\openone$.
The quasiclassical Green function obeys the Eilenberger equation [77],
@@
and is supplemented by the normalization condition, $\check{g}^{2}=-\pi^{2}$. $\check{\Delta}$ is the superconducting order parameter matrix. For spin-triplet superconductors with the $d$-vector, $\bm{d}(\bm{k})$, this order parameter matrix is given by [78, 76],
@@
where ${\bm{\sigma}}=(\sigma_{x},\sigma_{y},\sigma_{z})$ is the vector of the Pauli matrices in the spin space. $\check{\sigma}_{\rm imp}$ in Eq. (10) represents the impurity self-energy,
@@
As mentioned before, we consider short-range charge (scalar) impurities with the $\delta$-function type potential, $V_{\rm imp}(\bm{x})=\sum_{\bm{R}_{\rm imp}}V_{\rm imp}\delta(\bm{x}-\bm{R}_{\rm imp})$, where $\bm{R}_{\rm imp}$ is an impurity site and $V_{\rm imp}$ is the potential strength. The multiple scattering events are essential for the transverse transport and thus we compute the impurity self-energy with the self-consistent $T$-matrix approximation. Assuming the random distribution of impurities and taking the impurity average, we obtain the self-consistent $T$-matrix equation [26, 27],
@@
$\check{t}_{\rm imp}$ $=$ $V_{\rm imp}+N(\epsilon_{\rm F})V_{\rm imp}\braket{\check{g}}_{{\rm FS}}\check{t}_{\rm imp}.$ (12b)
where $n_{\rm imp}$ is the impurity density and $N(\epsilon_{\rm F})$ is the DOS at the Fermi level, $\epsilon_{\rm F}$, in the normal state. The bracket $\braket{\cdots}_{\rm FS}$ represents the normalized Fermi surface average, $\braket{1}_{\rm FS}=1$.
The impurity self-energy is independent of the Fermi momentum due to the short-range nature of the impurity potential. Using the scattering rate $\Gamma_{\rm imp}=n_{\rm imp}/\pi N(\epsilon_{\rm F})$ and the scattering phase-shift $\cot\delta=-1/\pi N(\epsilon_{\rm F})V_{\rm imp}$ in the normal state to parameterize the impurity scattering, we recast the self-consistent $T$-matrix equation as
$\check{\sigma}_{\rm imp}=-\left[\cot\delta+\left\langle\frac{\check{g}}{\pi}\right\rangle_{\rm FS}\right]^{-1}\Gamma_{\rm imp}$ (13)
The limit $\delta\to 0\;(\delta\to\frac{\pi}{2})$ corresponds to the Born (unitarity) limit with the weak (strong) impurity potential.
@@ IV.2 Response to Temperature Gradient
The quasiclassical formalism can treat the response to the temperature gradient [80, 81]. To involve the temperature gradient in the quasiclassical theory, we consider a local equilibrium, $T=T({\bm{x}})$, and expand the spatial gradient in Eq. (10) as $\bm{\nabla}=\bm{\nabla}T\frac{\partial}{\partial T}$,
@@
The thermal current is given by the Keldysh component of the Green function,
${\bm{J}}_{Q}$ $=$ $N(\epsilon_{\rm F})\int\frac{d\epsilon}{4\pi i}\left\langle\frac{1}{4}{\rm Tr}\left[\epsilon{\bm{v}_{\rm F}}\underline{g}^{\rm K}\right]\right\rangle_{\rm FS}\,,$ (15)
which we compute in the linear response theory, and then obtain the thermal conductivity tensor,
@@
As shown in Sec. III, we only obtain the extrinsic contribution to the THC in this framework. [10] Hence, we add the superscript “ext” in Eq. (16) to represent the extrinsic (impurity-induced) contribution.
Once the quasiclassical limit is taken, the anomalous velocity, the effective Lorentz force in the momentum space due to Berry curvature, and side-jump effects drop out from the equations for the integrated Green function [82, 83]. As shown in Sec. III, these geometric phase effects appear at first order in $(k_{\rm F}\xi_{0})^{-1}$, and accounting for them requires gradient expansion beyond the standard quasiclassical theory. [10] Consequently, when we discuss the extrinsic ATHE with the Eilenberger equation, it originates from the skew-like scattering due to chiral Cooper pairs. When we compare the extrinsic ATHE with the intrinsic one, we refer to the low temperature formula [Eq. (7)].
@@ IV.3 Quasiclassical Green function
With the quasiclassical transport theory, we derive the nonequilibrium Green function, which includes a linear response to the temperature gradient. Everywhere below the equilibrium functions are denoted as $\check{x}_{\rm eq}\;(x=g,\;\Delta,\sigma_{\rm imp})$ and their linear deviation from the equilibrium are labeled as $\delta\check{x}\;(x=g,\;\Delta,\sigma_{\rm imp})$.
IV.3.1 Equilibrium Green function
In the absence of perturbations, the equilibrium Green function, $\check{g}_{\rm eq}$, is obtained from Eq. (10) with the impurity self-energy given in Eq. (13). Because the Fermi surface average for the order parameter matrix vanishes, $\braket{\underline{\Delta}(\bm{k}_{\rm F})}_{\rm FS}=0$ due to the odd in momentum gap function, the impurity self-energy is diagonal in the Nambu space. The equation is then easily solved to obtain
$\underline{g}_{\rm eq}^{\rm X}$ $=$ $-\pi\frac{\underline{M}^{\rm X}}{D^{\rm X}}\;\;\;{\rm for\;X=R,\;A},$ (17)
@@ IV.3.1 Equilibrium Green function
IV.3.2 Nonequilibrium Green function
The nonequilibrium Green function, $\delta\check{g}$, describing the linear response to the temperature gradient obeys the Eilenberger equation,
$\left[\epsilon\check{\tau}_{z}-\check{\Delta}_{\rm eq}-\check{\sigma}_{\rm imp,eq},\delta\check{g}\right]-\left[\delta\check{\Delta}+\delta\check{\sigma}_{\rm imp},\check{g}_{\rm eq}\right]$
@@ IV.3.2 Nonequilibrium Green function
It is straightforward to solve this equation for the retarded and the advanced Green components. Using the normalization, $\{\underline{g}^{\rm X}_{\rm eq},\delta\underline{g}^{\rm X}\}=0\;({\rm X=R,A})$, we obtain,
$\delta\underline{g}^{\rm X}=\frac{\underline{g}^{\rm X}_{\rm eq}}{2\pi D^{\rm X}}\left(\left[\delta\underline{\sigma}^{\rm X}_{\rm imp},\underline{g}^{\rm X}_{\rm eq}\right]-(i{\bm{v}_{\rm F}}\cdot{\bm{\nabla}}T)\frac{\partial}{\partial T}\underline{g}^{\rm X}_{\rm eq}\right).$ (20)
The second term in Eq. (20) only depends on the temperature variations of the gap function, which is negligible at low temperatures. It can be neglected on even more general grounds because it is traceless, and hence does not contribute to the thermal transport. Indeed, it is easy to see that ${\rm Tr}(\underline{g}^{\rm X}_{\rm eq}\frac{\partial}{\partial T}\underline{g}^{\rm X}_{\rm eq})=2{\rm Tr}\left[\frac{\partial}{\partial T}(\underline{g}^{{\rm X}}_{\rm eq})^{2}\right]=0$.
For the Keldysh component, it is convenient to define the anomalous Keldysh Green function, $\delta\underline{g}^{a}$, and the anomalous Keldysh impurity self-energy, $\delta\underline{\sigma}_{\rm imp}^{a}$,
$\delta\underline{g}^{a}$ $=$ $\delta\underline{g}^{\rm K}-\left(\delta\underline{g}^{\rm R}-\delta\underline{g}^{\rm A}\right)\tanh\left(\frac{\epsilon}{2T}\right),$ (21)
$\delta\underline{\sigma}_{\rm imp}^{a}$ $=$ $\delta\underline{\sigma}_{\rm imp}^{\rm K}-\left(\delta\underline{\sigma}_{\rm imp}^{\rm R}-\delta\underline{\sigma}_{\rm imp}^{\rm A}\right)\tanh\left(\frac{\epsilon}{2T}\right).$ (22)
The second term in Eq. (21) describes the change in the spectral function, $\underline{g}^{\rm R}-\underline{g}^{\rm A}$ while maintaining the distribution in equilibrium. The first term in this equation accounts for the nonequilibrium distribution function and is essential for evaluating the thermal transport. This separation of the nonequilibrium Keldysh functions allows us to solve the transport equation for the Keldysh component. We obtain,
@@
where we defined the retarded function
@@
with the trace of the equilibrium self-energy $\sigma_{\rm imp,eq0}^{\rm X}={\rm Tr}(\sigma_{\rm imp,eq}^{\rm X})\;{\rm(X=R,A)}$. Note that $\delta\underline{g}^{a}_{\rm vc}$ involves the anomalous Keldysh impurity self-energy, which corresponds to the vertex correction in the diagrammatic calculations. We thus refer to $\delta\underline{g}^{a}_{\rm ns}$ as a non-selfconsistent contribution and $\delta\underline{g}^{a}_{\rm vc}$ as a vertex correction contribution, respectively.
@@ V Thermal conductivity at low temperature
As is seen above, the anomalous Keldysh Green function is proportional to the derivative of the Fermi distribution function, $1/\cosh^{2}(\epsilon/2T)$. This factor introduces the frequency cut-off $\epsilon\sim T$, which becomes small at low temperatures, justifying the expansion of the Green function in $\epsilon$ [22]. We utilize this low temperature expansion to determine the low temperature behavior of the extrinsic ATHE in WSCs.
Figure 2: The quasiparticle DOS for the WSCs with the $E_{1u}$ chiral pairings for the several values of the scattering phase-shift. In this calculation, we set the scattering rate as $\Gamma_{\rm imp}=0.04\pi T_{c,\rm clean}$ ($T_{c,\rm clean}$ is a critical temperature in clean systems) and the scattering phase-shift as $\delta=\frac{\pi}{12}$(red lines), $\frac{\pi}{6}$(blue lines), $\frac{\pi}{4}$(green lines), $\frac{\pi}{3}$(orange lines), $\frac{5\pi}{12}$(yellow lines), $\frac{\pi}{2}$(sky blue).
Figure 3: $\hat{k}_{z}$-dependence of the functions $\alpha_{0}$ and $\alpha_{1}$ in the WSCs with the $E_{1u}$ chiral pairings [see Eqs. (25)-(26)]. The red and green curves show the chiral $p$-wave and $f$-wave chiral pairings, respectively. We set the scattering rate $\Gamma_{\rm imp}=0.04\pi T_{c,\rm clean}$ and the phase-shift $\delta=\frac{\pi}{6}$.
Figure 4: Effect of the scattering phase shift. (a) The scattering phase-shift dependence of the extrinsic ATHE in WSCs at the low temperature. (b) The phase-shift dependence of the zero-energy DOS. In the panel (a), the dotted line plots $-\kappa_{xy}^{\rm int}$ with $0.011/(k_{\rm F}\xi_{0})=$ 0.01, where the minus sign of $-\kappa_{xy}^{\rm int}$ is due to compare the relative magnitude with $\kappa_{xy}^{\rm ext}$. In these calculations, we set the scattering rate as $\Gamma_{\rm imp}=0.04\pi T_{c,{\rm clean}}$ and the temperature as $T=0.02T_{c,{\rm clean}}$.
Figure 5: Temperature dependence of the extrinsic ATHE in the WSCs with the $E_{1u}$ chiral pairings. In all panels, we set the scattering rate as $\Gamma=0.04\pi T_{c}$ and the scattering phase-shift as $\delta=\frac{\pi}{12}$(red lines), $\frac{\pi}{6}$(blue lines), $\frac{\pi}{4}$(Green lines), $\frac{\pi}{3}$(orange lines), $\frac{5\pi}{12}$(yellow lines). The dotted lines plot $-\kappa_{xy}^{\rm int}$ using the low temperature formula with $0.011/(k_{\rm F}\xi_{0})=$ 0.01, where the minus sign of $-\kappa_{xy}^{\rm int}$ is due to compare the relative magnitude with $\kappa_{xy}^{\rm ext}$.
Following the procedure outlined in Ref. [22], we obtain the low temperature expansion for the thermal conductivity in WSCs as [84],
@@
with $i\gamma\equiv-\frac{1}{4}{\rm Tr}\left[\underline{\tau}_{z}\underline{\sigma}_{\rm imp,eq}^{\rm R}(\epsilon=0)\right]$. In Appendix B, we derive the complete expression for the thermal conductivity at the low temperature, including higher-order terms in the scattering rate, $\Gamma_{\rm imp}$. Note that the extrinsic contribution has the opposite sign of the intrinsic contribution when the impurity potential is attractive for electrons $(0<\delta<\pi/2)$ [22]. In Eqs. (25)-(26), the dimensionless factors $X$ and $Y$ are defined as,
@@
In Fig. 2, we plot the quasiparticle DOS in the $E_{1u}$ chiral $p$-wave and $f$-wave states. The Fermi surface function $\alpha_{n}(\hat{k}_{z})$ $(n\in\mathbb{Z})$, is defined as
@@
As shown in Fig. 3 (a), $\alpha_{0}(\hat{k}_{z})$ is sharply peaked at the positions of the line-nodes on the Fermi surface, because the denominator in Eq. (29) is small in these regions. Hence, as may be expected, low-energy quasiparticles near nodal lines enhance the longitudinal thermal transport. On the other hand, $\alpha_{1}(\hat{k}_{z})$ in Fig. 3(b) changes its sign across the line node due to $\eta(\hat{k}_{z})$ in the numerator of Eq. (29) (see also Fig. 1). As we show below, this sign change at line-nodes significantly affects the extrinsic ATHE.

### v2

@@ IV Quasiclassical transport theory
IV.1 Quasiclassical transport (Eilenberger) theory
The quasiclassical transport (Eilenberger) theory describes superconductors in the limit $(k_{\rm F}\xi_{0})^{-1}\ll 1$. In this limit, the normal state DOS can be taken to be energy-independent over the range where the Gor’kov Green function is peaked [77]. Integrating the Gor’kov Green function over the band kinetic energy, $\xi_{\bm{k}}={\bm{k}}^{2}/2m-\epsilon_{\rm F}$, where $\epsilon_{\rm F}$ is the Fermi energy, we define the quasiclassical Green function (QGF) as,
@@ IV.1 Quasiclassical transport (Eilenberger) theory
The QGF is defined at the Fermi surface and is the central object of the quasiclassical transport theory [77, 78]. The superscript $\rm X=R,A,K$ in Eq. (9) represents the retarded, advanced, and Keldysh matrix elements, respectively. $\check{\tau}_{\mu}\;(\mu=x,y,z)$ are the Pauli matrices in the Nambu (particle-hole) space. Throughout this paper, we denote $\check{A}$ as a $8\times 8$ matrix in the Keldysh space and $\underline{A}$ as a $4\times 4$ Nambu matrix. If a matrix $\underline{A}\;(\check{A})$ is defined in the Nambu (Keldysh) space, the corresponding matrix $\check{A}\;(\underline{A})$ in the Keldysh (Nambu) space is defined as $𝟙\check{A}=\underline{A}\otimes\openone$.
The QGF obeys the Eilenberger equation [77, 78],
@@
and is supplemented by the normalization condition,
$\check{g}^{2}=-\pi^{2}.$ (11)
This condition arises from the fact that the square of the QGF also satisfies the Eilenberger equation [77, 78]. $\check{\Delta}$ is the superconducting order parameter matrix. For spin-triplet superconductors with the $d$-vector, $\bm{d}(\bm{k})$, this order parameter matrix is written as [78, 76],
@@
where ${\bm{\sigma}}=(\sigma_{x},\sigma_{y},\sigma_{z})$ is the vector of the Pauli matrices in the spin space. $\check{\sigma}_{\rm imp}$ in Eq. (10) represents the impurity self-energy, which is expressed as,
@@
In this paper, we consider nonmagnetic impurities with the short-range $\delta$-function potential, $V_{\rm imp}(\bm{x})=\sum_{\bm{R}_{\rm imp}}V_{\rm imp}\delta(\bm{x}-\bm{R}_{\rm imp})$, where $\bm{R}_{\rm imp}$ is an impurity site and $V_{\rm imp}$ is the impurity potential strength. The multiple scattering events are essential for the transverse transport and thus we compute the impurity self-energy with the self-consistent $T$-matrix approximation [22, 23, 24]. Assuming the random distribution of impurities and taking the impurity average, we obtain the self-consistent $T$-matrix equation [26, 27],
@@
$\check{t}_{\rm imp}$ $=$ $V_{\rm imp}+N(\epsilon_{\rm F})V_{\rm imp}\braket{\check{g}}_{{\rm FS}}\check{t}_{\rm imp},$ (13b)
where $n_{\rm imp}$ is the impurity density and $N(\epsilon_{\rm F})$ is the DOS at the Fermi level in the normal state. The bracket $\braket{\cdots}_{\rm FS}$ represents the normalized Fermi surface average, $\braket{1}_{\rm FS}=1$. The impurity self-energy is independent of the Fermi momentum because we neglect the size of impurities.
We use the scattering rate $\Gamma_{\rm imp}=n_{\rm imp}/\pi N(\epsilon_{\rm F})$ and the scattering phase-shift $\cot\delta=-1/\pi N(\epsilon_{\rm F})V_{\rm imp}$ in the normal state to parameterize disorder in our calculation. Using these normal state quantities, we recast the self-consistent $T$-matrix equation as
$\check{\sigma}_{\rm imp}=-\left[\cot\delta+\left\langle\frac{\check{g}}{\pi}\right\rangle_{\rm FS}\right]^{-1}\Gamma_{\rm imp}.$ (14)
The limit $\delta\to 0\;(|\delta|\to\pi/2)$ corresponds to the Born (unitarity) limit, which represents the weak (strong) impurity potential limit, $V_{\rm imp}/\epsilon_{\rm F}\to 0$ ($|V_{\rm imp}/\epsilon_{\rm F}|\to\infty$).
@@ IV.2 Response to Temperature Gradient
The quasiclassical transport theory describes thermal responses [80, 81]. To include temperature gradients in the Eilenberger equation, we assume a local equilibrium, $T=T({\bm{x}})$, and expand the spatial gradient as $\bm{\nabla}=\bm{\nabla}T\frac{\partial}{\partial T}$. We then obtain,
@@
The thermal current is calculated from the Keldysh QGF as
${\bm{J}}_{Q}$ $=$ $N(\epsilon_{\rm F})\int\frac{d\epsilon}{4\pi i}\left\langle\frac{1}{4}{\rm Tr}\left[\epsilon{\bm{v}_{\rm F}}\underline{g}^{\rm K}\right]\right\rangle_{\rm FS}\,.$ (16)
Once we compute the thermal current within the linear response theory, we define the thermal conductivity tensor,
@@
As discussed above in Sec. III, in this framework, we only obtain the extrinsic contribution to the THC [10]. Hence, we add the superscript “ext” in Eq. (17) to represent the extrinsic (impurity-induced) contribution.
Once the quasiclassical limit is taken, the Berry curvature effects, including the anomalous velocity (effective Lorentz force in momentum space) and the side-jumps, drop out from the transport equation for the Green function [82, 83]. As shown in Sec. III, the Berry curvature effects appear from the first order in $(k_{\rm F}\xi_{0})^{-1}$, and the gradient expansion to higher orders is required to account for these effects [10]. Consequently, when we discuss the extrinsic ATHE using the Eilenberger equation, the ATHE originates from the skew scattering due to chiral Cooper pairs. When we compare the extrinsic ATHE with the intrinsic one, we refer to the low-temperature formula [Eq. (7)].
@@ IV.3 Quasiclassical Green function
Using the quasiclassical transport theory, we derive the QGF, which includes a linear response to the temperature gradient. Everywhere below the equilibrium functions are denoted as $\check{x}_{\rm eq}\;(x=g,\;\Delta,\sigma_{\rm imp})$ and their linear deviation from the equilibrium are labeled as $\delta\check{x}\;(x=g,\;\Delta,\sigma_{\rm imp})$.
IV.3.1 Equilibrium quasiclassical Green function
In global thermal equilibrium, the equilibrium QGF, $\check{g}_{\rm eq}$, is the solution of Eq. (10) without the gradient term. Because the Fermi surface average for the order parameter matrix vanishes in $p$- and $f$-wave superconductors, $\braket{\underline{\Delta}(\bm{k}_{\rm F})}_{\rm FS}=0$, the equilibrium impurity self-energy is diagonal in the Nambu space. Then, the Eilenberger equation can be easily solved using the normalization condition to give,
$\underline{g}_{\rm eq}^{\rm R,A}$ $=$ $-\pi\frac{\underline{M}^{\rm R,A}}{D^{\rm R,A}},$ (18)
@@ IV.3.1 Equilibrium quasiclassical Green function
IV.3.2 Nonequilibrium quasiclassical Green function
The nonequilibrium QGF, $\delta\check{g}$, which describes the linear response to the temperature gradient, obeys the Eilenberger equation,
$\left[\epsilon\check{\tau}_{z}-\check{\Delta}_{\rm eq}-\check{\sigma}_{\rm imp,eq},\delta\check{g}\right]-\left[\delta\check{\sigma}_{\rm imp},\check{g}_{\rm eq}\right]$
@@ IV.3.2 Nonequilibrium quasiclassical Green function
We can straightforwardly solve Eq. (IV.3.2) for the retarded and the advanced QGFs. Using the normalization condition $\{\underline{g}^{\rm R,A}_{\rm eq},\delta\underline{g}^{\rm R,A}\}=0$, we obtain,
$\delta\underline{g}^{\rm R,A}=\frac{\underline{g}^{\rm R,A}_{\rm eq}}{2\pi D^{\rm R,A}}\left(\left[\delta\underline{\sigma}^{\rm R,A}_{\rm imp},\underline{g}^{\rm R,A}_{\rm eq}\right]-(i{\bm{v}_{\rm F}}\cdot{\bm{\nabla}}T)\frac{\partial}{\partial T}\underline{g}^{\rm R,A}_{\rm eq}\right).$
The equilibrium retarded and advanced QGFs depend on the temperature only through the gap function, which makes the second term in Eq. (IV.3.2) negligible at low temperatures. That term is traceless, as can be easily verified using $(\underline{g}^{\rm R,A}_{\rm eq})^{2}=-\pi^{2}$ [22, 80],
${\rm Tr}\left[\underline{g}^{\rm R,A}_{\rm eq}\frac{\partial}{\partial T}\underline{g}^{\rm R,A}_{\rm eq}\right]$ $=$ $0\displaystyle 2{\rm Tr}\left[\frac{\partial}{\partial T}(\underline{g}^{{\rm R,A}}_{\rm eq})^{2}\right]=$ 0 . (22)
For the Keldysh component, it is convenient to define the anomalous Keldysh QGF, $\delta\underline{g}^{a}$, and the anomalous Keldysh impurity self-energy, $\delta\underline{\sigma}_{\rm imp}^{a}$,
$\delta\underline{g}^{\rm K}$ $=$ $\left(\delta\underline{g}^{\rm R}-\delta\underline{g}^{\rm A}\right)\tanh\left(\frac{\epsilon}{2T}\right)+\delta\underline{g}^{a},$ (23)
$\delta\underline{\sigma}_{\rm imp}^{K}$ $=$ $\left(\delta\underline{\sigma}_{\rm imp}^{\rm R}-\delta\underline{\sigma}_{\rm imp}^{\rm A}\right)\tanh\left(\frac{\epsilon}{2T}\right)+\delta\underline{\sigma}_{\rm imp}^{\rm a}.$ (24)
The first term in Eq. (23) describes the change in the spectral function, which is proportional to $\underline{g}^{\rm R}-\underline{g}^{\rm A}$ while maintaining the distribution in equilibrium. This term does not contribute to the thermal conductivity [80]. The second term in Eq. (23) accounts for the change in the distribution function from equilibrium and is essential for evaluating thermal transport properties.
This separation of the nonequilibrium Keldysh QGF simplifies the transport equation and allows us to solve the Eilenberger equation for the Keldysh component. Using the constraint $\underline{g}^{{\rm R}}_{\rm eq}\delta\underline{g}^{{a}}+\delta\underline{g}^{{a}}\underline{g}^{{\rm A}}_{\rm eq}=0$ from the normalization condition, we obtain,
@@
where we defined the retarded function as,
@@
with the trace of the equilibrium self-energy $\sigma_{\rm imp,eq0}^{\rm R,A}={\rm Tr}(\sigma_{\rm imp,eq}^{\rm R,A})$. Note that $\delta\underline{g}^{a}_{\rm vc}$ involves $\delta\underline{\sigma}_{\rm imp}^{a}$, which includes the vertex correction in the diagrammatic calculations. We thus refer to $\delta\underline{g}^{a}_{\rm ns}$ as a non-selfconsistent contribution and $\delta\underline{g}^{a}_{\rm vc}$ as a vertex correction contribution, respectively.
@@ V Thermal conductivity at low temperature
As shown in Eq. (23), the anomalous Keldysh QGF is proportional to the derivative of the Fermi distribution function, $\partial_{\epsilon}f(\epsilon)\propto 1/\cosh^{2}(\epsilon/2T)$. This factor peaks at $\epsilon=0$ and introduces the frequency cut-off $\epsilon_{\rm cut}\sim T$. This cut-off frequency becomes small at low temperatures, justifying the expansion of the QGF in $\epsilon$ [22]. We perform this low-temperature expansion to obtain the low-temperature formula for the extrinsic ATHE.
Figure 2: Quasiparticle DOS in (a) the chiral $p$-wave and (b) the chiral $f$-wave states for the several values of the scattering phase-shift. In this calculation, we set the scattering rate $\Gamma_{\rm imp}=0.04\pi T_{c,\rm clean}$ ($T_{c,\rm clean}$ is a critical temperature in clean systems) and the scattering phase-shift $\delta=\frac{\pi}{12}$(red curves), $\frac{\pi}{6}$(blue curves), $\frac{\pi}{4}$(green curves), $\frac{\pi}{3}$(orange curves), $\frac{5\pi}{12}$(yellow curves), $\frac{\pi}{2}$(sky blue curves).
Following the procedure outlined in Ref. [22], we obtain the low-temperature formula for the thermal conductivity as [84],
$\frac{\kappa^{\rm ext}_{yy}}{N(\epsilon_{\rm F})v_{\rm F}^{2}}=$ $\frac{\pi^{2}T}{6}\gamma^{2}\braket{\alpha_{0}(\hat{k}_{z})}_{{\rm FS}}$
@@
$\frac{\kappa^{\rm ext}_{xy}}{N(\epsilon_{\rm F})v_{\rm F}^{2}}=$ $-\frac{\pi^{2}\Gamma_{\rm imp}\gamma^{2}|\Delta_{\rm eq}|^{2}T}{3}X\braket{\alpha_{1}(\hat{k}_{z})}_{{\rm FS}}^{2}$
@@
with $i\gamma\equiv-\frac{1}{4}{\rm Tr}\left[\underline{\tau}_{z}\underline{\sigma}_{\rm imp,eq}^{\rm R}(\epsilon=0)\right]$. In Appendix B, we derive the complete expression for the thermal conductivity at the low temperature, including higher-order terms in the scattering rate, $\Gamma_{\rm imp}$. In Eqs. (27) and (28), the dimensionless factors $X$ and $Y$ are defined as,
@@
Additionally, the Fermi surface functions $\alpha_{n}(\hat{k}_{z})$ $(n\in\mathbb{Z})$ are defined as
@@
Figure 3: Fermi surface functions (a) $\alpha_{0}(\hat{k}_{z})$ and (b) $\alpha_{1}(\hat{k}_{z})$ in the chiral $p$-wave and the $f$-wave states. In all panels, the red and green curves represent the calculated results for the chiral $p$-wave and $f$-wave states, respectively. The positions of the line-nodes in the chiral $f$-wave state ($\hat{k}_{z}=\pm 1/\sqrt{5}$) are indicated by the vertical dashed lines. In these calculations, we set the scattering rate $\Gamma_{\rm imp}=0.04\pi T_{c,\rm clean}$ and the phase-shift $\delta=\frac{\pi}{6}$.
The factors $X$ and $Y$ stem from the vertex correction contributions to the nonequilibrium Keldysh QGF. These factors are closely related to impurity-bound states in superconductors, and we briefly review the common physics of these states before discussing the thermal transport properties [28, 29]. In unconventional superconductors with a momentum-dependent gap function, multiple Andreev scattering at impurity sites creates quasiparticle-bound states. In our formulation, such impurity effects are included in the impurity self-energy under the self-consistent $T$-matrix approximation. The impurity self-energy introduces new poles in the equilibrium QGF. These new poles occur at both positive (electron-like) and negative (hole-like) energies due to the particle-hole symmetric spectrum. The impurity level is positioned at $\epsilon=0$ in the unitarity limit ($|\delta|=\pi/2$), and moves to a finite energy when the phase-shift deviates from this limit. For the weak impurity potential ($|\delta|\ll\pi/2$), the spectral weight for the bound states merges with the coherence peaks in the DOS at $\epsilon=\pm|\Delta_{\rm eq}|$. For a finite impurity density, the impurity-bound states broaden into impurity bands, centered at the energy of single impurity resonant states, with a finite bandwidth determined by the details of impurity scattering.
Figure 2 shows the DOS for the several values of $\delta$, illustrating the phase-shift dependence explained above. When the impurity potential is weak (such as the red, green and blue curves in Fig. 2 (a-b)), the low-energy DOS ($|\epsilon|=|\Delta_{\rm eq}|$) varies with the energy as $N_{s}(\epsilon)\propto\epsilon^{2}$ for the chiral $p$ wave state, and $N_{s}(\epsilon)\propto|\epsilon|$ for the chiral $f$ wave state, reflecting the dimensionality of the nodal regions in the gap structure [76]. As $\delta$ increases, the spectral weight in the DOS is partially transferred from the coherence peaks at $|\epsilon|=|\Delta_{\rm eq}|$ into the mid-gap region $|\epsilon|<|\Delta_{\rm eq}|$. As the scattering phase-shift approaches the unitarity limit, even in the dilute impurity limit that we consider, the transferred spectral weight shifts close to the Fermi energy, with the tails of the impurity bands touching $\epsilon=0$ [see the orange, yellow, and sky blue curves in Fig. 2 (a-b)].
Figure 4: Scattering phase-shift dependence of (a) the impurity contribution to the THC, and (b) the zero-energy DOS in the $E_{1u}$ chiral $p$-wave and $f$-wave states at the low temperature. In all panels, The red and green curves represent the calculated results for the chiral $p$-wave and $f$-wave states. In panel (a), the dotted line plots $-\kappa_{xy}^{\rm int}T_{c,{\rm clean}}/(N(\epsilon_{\rm F})v_{\rm F}^{2}T)$ with $(k_{\rm F}\xi_{0})^{-1}=0.01$, where the minus sign is due to compare the relative magnitude with $\kappa_{xy}^{\rm ext}$. In panel (b), the dashed vertical lines indicate the peak positions of the THC as a function of the scattering phase-shift for the chiral $p$-wave state (red line) and the chiral $f$-wave state (green line). In these calculations, we set the scattering rate $\Gamma_{\rm imp}=0.04\pi T_{c,{\rm clean}}$ and the temperature $T=0.02T_{c,{\rm clean}}$.
Figure 5: Temperature dependence of (a) the impurity contribution to the THC and (b) the longitudinal thermal conductivity in the $E_{1u}$ chiral $p$-wave and $f$-wave states. In panel (a), the dotted lines plot $-\kappa_{xy}^{\rm int}/(N(\epsilon_{\rm F})v_{\rm F}^{2})$ using the low-temperature formula with $(k_{\rm F}\xi_{0})^{-1}=0.01$, where the minus sign is due to comparing the relative magnitude with $\kappa_{xy}^{\rm ext}$. In all panels, we set the scattering rate $\Gamma_{\rm imp}=0.04\pi T_{c,{\rm clean}}$ and the scattering phase-shift $\delta=\frac{\pi}{12}$ (red curves), $\frac{\pi}{6}$ (blue curves), $\frac{\pi}{4}$ (green curves), $\frac{\pi}{3}$ (orange curves), $\frac{5\pi}{12}$ (yellow curves).
We now focus on the thermal transport coefficients at low temperatures, including their sign and magnitude. The impurity contribution to the THC is proportional to the factor $X$, which has a positive (negative) value when the impurity potential is attractive (repulsive) for electrons. Referring to the low-temperature formula for the intrinsic contribution [Eq. (7)], we find that the intrinsic and impurity contributions to the THC have opposite signs for the attractive impurity potential ($20<\delta<\pi/$ 2[22].
The first term in Eq. (27) is proportional to $\braket{\alpha_{0}(\hat{k}_{z})}_{{\rm FS}}$. As shown in Fig. 3 (a), $\alpha_{0}(\hat{k}_{z})$ is positive and sharply peaks at the positions of the line-nodes on the Fermi surface, making its Fermi surface average large. Hence, line-nodal excitations always enhance the longitudinal thermal transport [see also Fig. 5 (b)]. In contrast, whether line-nodal excitations enhance the extrinsic ATHE depends on the details of the order parameters [see also Sec. VII]. The impurity contribution to the THC is proportional to the square of the Fermi surface average of $\alpha_{1}(\hat{k}_{z})$. $\alpha_{1}(\hat{k}_{z})$ changes sign across line-nodes due to $\eta(\hat{k}_{z})$ in the numerator, which reduces its average over the Fermi surface. Figure 3 (b) clearly shows that in the chiral $f$-wave state, the two horizontal line-nodes involve such sign changes in $\alpha_{1}(\hat{k}_{z})$. As we show in Sec. VII, this sign change at line nodes significantly affects the extrinsic ATHE.